Optical equipment is purchased on December 1, 2018 for $7,500.00 The useful life for
this fixed asset is five years, with a salvage value of $1,500.00. A new computer is also
purchased on December 1, 2018 for $1,800.00. The useful life for this fixed asset is three
years with a salvage value of $50.00.
What will be the general journal entry to record depreciation on December 31, 2018
by using the straight-line method?
Answer #1

Answer

  • Working for Optical Equipment

A

Cost

$              7,500.00

B

Residual Value

$              1,500.00

C=A - B

Depreciable base

$              6,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$              1,200.00

--No. of month asset used in 2018 = 1 month [December].
--Depreciation expense = 1200 x 1 month/12 months = $ 100

  • Working for Computer

A

Cost

$              1,800.00

B

Residual Value

$                    50.00

C=A - B

Depreciable base

$              1,750.00

D

Life [in years]

3

E=C/D

Annual SLM depreciation

$                  583.33

--No. of month asset used in 2018 = 1 month [December].
--Depreciation expense = 583.333 x 1 month/12 months = $ 48.61

  • Journal Entry

Date

Accounts title

Debit

Credit

31-Jan-18

Depreciation expense – Equipment [$ 100 + $ 48.61]

$            148.61

Accumulated Depreciation - Equipment

$        148.61

(depreciation expense recorded)